# The numeraire e-variable and reverse information projection

### v3

@@ 4 Finding the numeraire
Assume that ${\mathsf{Q}}\ll{\mathcal{P}}$. Let $X^{*}$ be an e-variable that is ${\mathsf{Q}}$-almost surely strictly positive and let ${\mathsf{P}}^{*}$ be given by $d{\mathsf{P}}^{*}/d{\mathsf{Q}}=1/X^{*}$. Then $X^{*}$ is a numeraire if and only if ${\mathsf{P}}^{*}$ belongs to the effective null ${\mathcal{P}}^{\circ\circ}$. Thus the numeraire is the only e-variable that is also a likelihood ratio between ${\mathsf{Q}}$ and some equivalent element of ${\mathcal{P}}^{\circ\circ}$. Finally, ${\mathbb{E}}_{{\mathsf{P}}^{*}}[X^{*}]=1$ and ${\mathsf{P}}^{*}$ is maximal in the sense that no other ${\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}$ equivalent to ${\mathsf{Q}}$ can satisfy ${\mathsf{P}}(A)\geq{\mathsf{P}}^{*}(A)$ for all $A$ with strict inequality for some $A$.
@@
Recall that all e-variables are ${\mathsf{Q}}$-almost surely finite thanks to Theorem 2.5 and the assumption that ${\mathsf{Q}}\ll{\mathcal{P}}$; in particular, ${\mathsf{P}}^{*}$ is equivalent to ${\mathsf{Q}}$. Now, if $X^{*}$ is a numeraire, then Lemma 3.1 yields that ${\mathsf{P}}^{*}\in{\mathcal{P}}^{\circ\circ}$. For the converse, the definition of ${\mathsf{P}}^{*}$ and the fact that it belongs to ${\mathcal{P}}^{\circ\circ}$ yield ${\mathbb{E}}_{\mathsf{Q}}[X/X^{*}]={\mathbb{E}}_{{\mathsf{P}}^{*}}[X]\leq 1$ for any e-variable $X$. This is the numeraire property. Finally, the definition of ${\mathsf{P}}^{*}$ immediately gives ${\mathbb{E}}_{{\mathsf{P}}^{*}}[X^{*}]=1$, and ${\mathsf{P}}^{*}$ must be maximal because if a ‘dominating’ equivalent ${\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}$ existed we would get the contradiction $1={\mathbb{E}}_{{\mathsf{P}}^{*}}[X^{*}]<{\mathbb{E}}_{\mathsf{P}}[X^{*}]\leq 1$. ∎
The following corollary applies in settings where the null hypothesis is ‘generated’, in the sense of (4.1) below, by a subset ${\mathcal{E}}_{0}$ of e-variables. Null hypotheses of this kind were studied by Casgrain et al. (2024) in a sequential setting. The corollary shows that once a candidate $X^{*}$ has been found, it is enough to check the numeraire property for the generating family ${\mathcal{E}}_{0}$, provided that it holds with equality for the e-variable which is identically one. The latter condition means that the RIPr ${\mathsf{P}}^{*}$ is a probability measure.
@@
Let ${\mathsf{P}}^{*}$ be given by $d{\mathsf{P}}^{*}/d{\mathsf{Q}}=1/X^{*}$. Then (4.2) says that ${\mathsf{P}}^{*}$ is a probability measure such that ${\mathbb{E}}_{{\mathsf{P}}^{*}}[X]\leq 1$ for all $X\in{\mathcal{E}}_{0}$. Thus by (4.1), ${\mathsf{P}}^{*}$ belongs to ${\mathcal{P}}$, hence to ${\mathcal{P}}^{\circ\circ}$. The result now follows from Theorem 4.1. ∎
@@
Next, we consider the dual approach of finding ${\mathsf{P}}^{*}$ directly, either by minimizing relative entropy or, if a candidate is available, checking that (ii) or (iii) of Theorem 3.4 are satisfied. The difficulty is that the domain is all of ${\mathcal{P}}^{\circ\circ}$, which can be hard to describe explicitly. Matters would become simpler if one could instead work with ${\mathcal{P}}$. We first show that this is not possible in general.
@@
This example shows that Theorem 3.4(ii) may no longer characterize the RIPr ${\mathsf{P}}^{*}$ uniquely if ${\mathcal{P}}^{\circ\circ}$ is replaced by ${\mathcal{P}}$. Let $\Omega=[0,1]$ with its Borel $\sigma$-algebra, and let ${\mathcal{P}}=\{\delta_{\omega}\colon\omega\in[0,1]\}$ be the set of all Dirac point masses. In this case the set of e-variables is trivial in the sense that a random variable $X$ is an e-variable if and only if $0\leq X(\omega)\leq 1$ for all $\omega\in[0,1]$. It follows that ${\mathcal{P}}^{\circ\circ}$ contains all sub-probabilities on $\Omega$. Now, let the alternative hypothesis ${\mathsf{Q}}$ be the uniform distribution on the unit interval. We then have the extreme situation that ${\mathcal{P}}$ does not contain any distribution absolutely continuous with respect to ${\mathsf{Q}}$, whereas ${\mathcal{P}}^{\circ\circ}$ contains all such distributions. In particular, Theorem 3.4(ii) would be vacuously true for any candidate ${\mathsf{P}}^{*}$ if ${\mathcal{P}}^{\circ\circ}$ were replaced by ${\mathcal{P}}$. In contrast, as currently stated, Theorem 3.4(ii) will only hold for ${\mathsf{P}}^{*}={\mathsf{Q}}$, which is the correct answer.
The situation improves in the presence of a reference measure $\mu$ with respect to which ${\mathsf{Q}}$ and all ${\mathsf{P}}\in{\mathcal{P}}$ are absolutely continuous. In this setting, if ${\mathsf{P}}_{n},{\mathsf{P}}$ are measures with densities $p_{n}=d{\mathsf{P}}_{n}/d\mu$ and $p=d{\mathsf{P}}/d\mu$ we say that ${\mathsf{P}}_{n}$ $\mu$-converges to ${\mathsf{P}}$ if $p_{n}$ converges to $p$ in probability under $\mu$. A set $C$ of absolutely continuous measures is called $\mu$-closed if it is closed with respect to $\mu$-convergence. It is called solid if it contains every (nonnegative) measure ${\mathsf{P}}^{\prime}$ that is dominated by some ${\mathsf{P}}\in C$ in the sense that ${\mathsf{P}}^{\prime}(A)\leq{\mathsf{P}}(A)$ for all $A$. The following result characterizes ${\mathcal{P}}^{\circ\circ}$ in terms of these concepts.
@@
Suppose all ${\mathsf{P}}\in{\mathcal{P}}$ are absolutely continuous with respect to a probability measure $\mu$. Then all elements of ${\mathcal{P}}^{\circ\circ}$ are also absolutely continuous with respect to $\mu$, and ${\mathcal{P}}^{\circ\circ}$ is the smallest $\mu$-closed convex solid set that contains ${\mathcal{P}}$.
@@
We identify ${\mathcal{P}}$ with its set of densities $C=\{p\in L^{0}_{+}(\mu)\colon p=d{\mathsf{P}}/d\mu,\ {\mathsf{P}}\in{\mathcal{P}}\}$. The polar of $C$ in $L^{0}_{+}(\mu)$ is $C^{\circ}=\{f\in L^{0}_{+}(\mu)\colon{\mathbb{E}}_{\mu}[pf]\leq 1\text{ for all }p\in C\}$, which is the set of (equivalence classes of) random variables that are $\mu$-almost surely equal to a finite e-variable. We claim that all elements of ${\mathcal{P}}^{\circ\circ}$ are absolutely continuous with respect to $\mu$, and that its set of densities is precisely the bipolar $C^{\circ\circ}=\{p\in L^{0}_{+}(\mu)\colon{\mathbb{E}}_{\mu}[pf]\leq 1\text{ for all }f\in C^{\circ}\}$. To check absolute continuity, let ${\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}$ and consider any event $A$ such that $\mu(A)=0$. Then $\infty\bm{1}_{A}$ is an e-variable, hence ${\mathbb{E}}_{\mathsf{P}}[\infty\bm{1}_{A}]\leq 1$, and thus ${\mathsf{P}}(A)=0$. This shows absolute continuity and lets us identify ${\mathcal{P}}^{\circ\circ}$ with its set of densities, namely, the set of $p\in L^{0}_{+}(\mu)$ such that ${\mathbb{E}}_{\mu}[pX]\leq 1$ for all finite e-variables $X$. Here we may restrict to finite e-variables because ${\mathbb{E}}_{\mu}[pX]\leq 1$ if and only if ${\mathbb{E}}_{\mu}[p(X\wedge n)]\leq 1$ for all $n\in{\mathbb{N}}$. The finite e-variables are, up to $\mu$-nullsets, precisely the elements of $C^{\circ}$, so we deduce that ${\mathcal{P}}^{\circ\circ}$ is indeed identified with $C^{\circ\circ}$. Finally, the bipolar theorem of Brannath and Schachermayer (1999, Theorem 1.3) states that $C^{\circ\circ}$ is the smallest $\mu$-closed convex solid set that contains $C$. This completes the proof. ∎
@@
If ${\mathcal{P}}$ is finite, then ${\mathcal{P}}^{\circ\circ}$ consists of all sub-probabilities that are set-wise dominated by an element of $\operatorname{conv}({\mathcal{P}})$. In particular, ${\mathcal{P}}^{\circ\circ}\cap M_{1}=\operatorname{conv}({\mathcal{P}})$.
@@
For finite ${\mathcal{P}}=\{{\mathsf{P}}_{1},\ldots,{\mathsf{P}}_{n}\}$ we may take $\mu=({\mathsf{P}}_{1}+\cdots+{\mathsf{P}}_{n})/n$ as reference measure. In this case $\operatorname{conv}({\mathcal{P}})$ is already $\mu$-closed and the form of ${\mathcal{P}}^{\circ\circ}$ follows from Lemma 4.5. ∎
Lemma 4.5 is powerful because it characterizes all elements of ${\mathcal{P}}^{\circ\circ}$. The following theorem makes use of this to simplify the conditions of Theorem 3.4. Lemma 4.5 also provides a way of checking that a particular measure belongs to ${\mathcal{P}}^{\circ\circ}$. For example, it shows that limits in probability of (densities of) measures in the convex hull of ${\mathcal{P}}$ must belong to ${\mathcal{P}}^{\circ\circ}$. The latter is however also easy to check directly from the definition of ${\mathcal{P}}^{\circ\circ}$ and Fatou’s lemma, and does not require the bipolar theorem used in the proof of Lemma 4.5.
If a reference measure $\mu$ is given, we identify any ${\mathsf{P}}\in{\mathcal{P}}$ with its density $p=d{\mathsf{P}}/d\mu$. Thanks to Lemma 4.5, we may similarly identify ${\mathcal{P}}^{\circ\circ}$ with its set of densities. For the sake of brevity we abuse notation and write, for example, $p\in{\mathcal{P}}$ even though, strictly speaking, ${\mathcal{P}}$ is a set of measures, not densities. Next, given ${\mathsf{Q}}$ with density $q$, the absolutely continuous part of any $p\in{\mathcal{P}}^{\circ\circ}$ has density $p^{a}=p\bm{1}_{\{q>0\}}$, and this is ${\mathsf{Q}}$-almost surely equal to $p$ itself. In view of these observations, the conditions (ii) and (iii) of Theorem 3.4 characterizing the RIPr state, respectively, that ${\mathbb{E}}_{\mathsf{Q}}[p/p^{*}]\leq 1$ and ${\mathbb{E}}_{\mathsf{Q}}[\log(p/p^{*})]\leq 0$, for all $p\in{\mathcal{P}}^{\circ\circ}$. Here $p^{*}=d{\mathsf{P}}^{*}/d\mu$. The following theorem simplifies these conditions by showing that, essentially, it is enough to check them for elements of ${\mathcal{P}}$ rather than ${\mathcal{P}}^{\circ\circ}$.
@@
Assume that ${\mathsf{Q}}\ll{\mathcal{P}}$ and that ${\mathsf{Q}}$ and all elements of ${\mathcal{P}}$ are absolutely continuous with respect to a probability measure $\mu$. Let $p^{*}\in{\mathcal{P}}^{\circ\circ}$ be equivalent to ${\mathsf{Q}}$. Then each of the conditions (i), (ii), and (iii) of Theorem 3.4 is equivalent to
@@
The reason for the asymmetry between Theorem 3.4(iii) and (4.4) is that the latter only involves $p$ from ${\mathcal{P}}$, not ${\mathcal{P}}^{\circ\circ}$. Since ${\mathcal{P}}$ may not contain $p^{*}$, nor be convex, we need to explicitly include convex combinations with $p^{*}$ in the numerator on the left-hand side of (4.4). Let us also emphasize that $t=1$ is excluded in (4.4). This sometimes makes the condition easier to check, for instance in the proof of Corollary 4.9 below.
@@
We know that (i), (ii), and (iii) of Theorem 3.4 are equivalent for elements of ${\mathcal{P}}^{\circ\circ}$ that are equivalent to ${\mathsf{Q}}$. To prove the theorem it is therefore enough to show that that Theorem 3.4(ii), (4.3), and (4.4) are equivalent.
We first show that Theorem 3.4(ii) is equivalent to (4.3). The former is clearly a stronger condition because it states that (4.3) holds with the larger set ${\mathcal{P}}^{\circ\circ}$ in place of ${\mathcal{P}}$. For the converse, suppose (4.3) holds. Let ${\mathcal{P}}^{\prime}$ be the set of all sub-probability densities dominated by some element of the convex hull of ${\mathcal{P}}$, and let ${\mathcal{P}}^{\prime\prime}$ be the closure in $\mu$-probability of ${\mathcal{P}}^{\prime}$. It is clear that (4.3) holds for all $p\in{\mathcal{P}}^{\prime}$ and then, by Fatou’s lemma, also for all $p\in{\mathcal{P}}^{\prime\prime}$. We claim that ${\mathcal{P}}^{\prime\prime}$ is a closed convex solid set that contains ${\mathcal{P}}$; it must then also contain ${\mathcal{P}}^{\circ\circ}$ since this is the smallest such set by Lemma 4.5. From this we conclude that Theorem 3.4(ii) holds. To see that ${\mathcal{P}}^{\prime\prime}$ is closed, convex, and solid, note that by definition, every $p^{\prime\prime}\in{\mathcal{P}}^{\prime\prime}$ is the limit in probability of some $p_{n}^{\prime}\in{\mathcal{P}}^{\prime}$, meaning that $p_{n}^{\prime}\leq p_{n}$ for some $p_{n}$ in the convex hull of ${\mathcal{P}}$. From this description convexity follows directly, closedness is obtained from a diagonal argument, and solidity follows by noting that any $\widetilde{p}\leq p^{\prime\prime}$ is the limit of $\widetilde{p}\wedge p_{n}^{\prime}$, which is still dominated by $p_{n}$ and thus belongs to ${\mathcal{P}}^{\prime}$.
@@
if ${\mathsf{Q}}$ is absolutely continuous with respect to $p$, and +$\infty$+ ∞ otherwise. The following two corollaries confirm, in particular, that the definition of the RIPr used in this paper, Definition 3.2, is consistent with existing definitions in the literature.
@@
If ${\mathcal{P}}$ is convex and $(p_{n})_{n\in{\mathbb{N}}}$ is a sequence in ${\mathcal{P}}$ with $H({\mathsf{Q}}\mid p_{n}\leadsto{\mathcal{P}})\to 0$, then $\log p_{n}$ converges to $\log p^{*}$ in $L^{1}({\mathsf{Q}})$, where $p^{*}$ is the RIPr.
@@
Theorem 4.7 now shows that $p^{*}$ is the RIPr, provided we can argue that $p^{*}$ belongs to ${\mathcal{P}}^{\circ\circ}$. This will use that $p^{*}=0$ on $\{q=0\}$. Specifically, note that ${\mathcal{P}}^{\circ\circ}$ contains $p_{n}$ and hence, being solid, also the smaller random variable $p_{n}\bm{1}_{\{q>0\}}$. Convergence in ${\mathsf{Q}}$-probability of $p_{n}$ to $p^{*}$ implies convergence in $\mu$-probability of $p_{n}\bm{1}_{\{q>0\}}$ to $p^{*}\bm{1}_{\{q>0\}}=p^{*}$. Since ${\mathcal{P}}^{\circ\circ}$ is closed in $\mu$-probability, it must contain $p^{*}$. This completes the proof. ∎

### v4

@@ 4 Finding the numeraire and RIPr
Assume that ${\mathsf{Q}}\ll{\mathcal{P}}$. Let $X^{*}$ be a ${\mathsf{Q}}$-almost surely strictly positive e-variable and let ${\mathsf{P}}^{*}$ be the measure given by $d{\mathsf{P}}^{*}/d{\mathsf{Q}}=1/X^{*}$. Then $X^{*}$ is a numeraire if and only if ${\mathsf{P}}^{*}$ belongs to the effective null ${\mathcal{P}_{\textnormal{eff}}}$. Thus the numeraire is the only e-variable that is also a likelihood ratio between ${\mathsf{Q}}$ and some equivalent element of ${\mathcal{P}_{\textnormal{eff}}}$. Finally, ${\mathbb{E}}_{{\mathsf{P}}^{*}}[X^{*}]=1$ and ${\mathsf{P}}^{*}$ is maximal in the sense that no other ${\mathsf{P}}\in{\mathcal{P}_{\textnormal{eff}}}$ equivalent to ${\mathsf{Q}}$ can satisfy ${\mathsf{P}}(A)\geq{\mathsf{P}}^{*}(A)$ for all $A$ with strict inequality for some $A$.
@@
Recall that all e-variables are ${\mathsf{Q}}$-almost surely finite thanks to Theorem 2.6 and the assumption that ${\mathsf{Q}}\ll{\mathcal{P}}$; in particular, ${\mathsf{P}}^{*}$ is equivalent to ${\mathsf{Q}}$. Now, if $X^{*}$ is a numeraire, then Lemma 3.3 yields that ${\mathsf{P}}^{*}\in{\mathcal{P}_{\textnormal{eff}}}$. For the converse, the definition of ${\mathsf{P}}^{*}$ and the fact that it belongs to ${\mathcal{P}_{\textnormal{eff}}}$ yield ${\mathbb{E}}_{\mathsf{Q}}[X/X^{*}]={\mathbb{E}}_{{\mathsf{P}}^{*}}[X]\leq 1$ for any e-variable $X$. This is the numeraire property. Finally, the definition of ${\mathsf{P}}^{*}$ immediately gives ${\mathbb{E}}_{{\mathsf{P}}^{*}}[X^{*}]=1$, and ${\mathsf{P}}^{*}$ must be maximal because if a ‘dominating’ equivalent ${\mathsf{P}}\in{\mathcal{P}_{\textnormal{eff}}}$ existed we would get the contradiction $1={\mathbb{E}}_{{\mathsf{P}}^{*}}[X^{*}]<{\mathbb{E}}_{\mathsf{P}}[X^{*}]\leq 1$. ∎
4.1 Generated null hypotheses
The following corollary of Theorem 4.1 applies in settings where the null hypothesis is ‘generated’, in the sense of (4.1) below, by a subset ${\mathcal{E}}_{0}$ of e-variables. Null hypotheses of this kind were studied by Casgrain et al. (2024) in a sequential setting. The corollary shows that once a candidate $X^{*}$ has been found, it is enough to check the numeraire property for the generating family ${\mathcal{E}}_{0}$, provided that it holds with equality for the e-variable which is identically one. The latter condition means that the RIPr ${\mathsf{P}}^{*}$ is a probability measure.
@@ 4.1 Generated null hypotheses
Let ${\mathsf{P}}^{*}$ be given by $d{\mathsf{P}}^{*}/d{\mathsf{Q}}=1/X^{*}$. Then (4.2) says that ${\mathsf{P}}^{*}$ is a probability measure such that ${\mathbb{E}}_{{\mathsf{P}}^{*}}[X]\leq 1$ for all $X\in{\mathcal{E}}_{0}$. Thus by (4.1), ${\mathsf{P}}^{*}$ belongs to ${\mathcal{P}}$, hence to ${\mathcal{P}_{\textnormal{eff}}}$. The result now follows from Theorem 4.1. ∎
@@
4.2 When a reference measure exists
We now consider the dual approach of finding ${\mathsf{P}}^{*}$ directly, either by minimizing relative entropy or, if a candidate is available, checking that (ii) or (iii) of Theorem 3.6 are satisfied. The difficulty is that the domain is all of ${\mathcal{P}_{\textnormal{eff}}}$, which can be hard to describe explicitly. Matters would become simpler if one could instead work directly with ${\mathcal{P}}$, but this is not possible in general. Here is one example of what can go wrong.
@@ 4.2 When a reference measure exists
This example shows that Theorem 3.6(ii) may no longer characterize the RIPr ${\mathsf{P}}^{*}$ uniquely if ${\mathcal{P}_{\textnormal{eff}}}$ is replaced by ${\mathcal{P}}$. Let $\Omega=[0,1]$ with its Borel $\sigma$-algebra, and let ${\mathcal{P}}=\{\delta_{\omega}\colon\omega\in[0,1]\}$ be the set of all Dirac point masses. In this case the set of e-variables is trivial in the sense that a random variable $X$ is an e-variable if and only if $0\leq X(\omega)\leq 1$ for all $\omega\in[0,1]$. It follows that ${\mathcal{P}_{\textnormal{eff}}}$ contains all sub-probabilities on $\Omega$. Now, let the alternative hypothesis ${\mathsf{Q}}$ be the uniform distribution on the unit interval. We then have the extreme situation that ${\mathcal{P}}$ does not contain any distribution absolutely continuous with respect to ${\mathsf{Q}}$, whereas ${\mathcal{P}_{\textnormal{eff}}}$ contains all such distributions. In particular, Theorem 3.6(ii) would be vacuously true for any candidate ${\mathsf{P}}^{*}$ if ${\mathcal{P}_{\textnormal{eff}}}$ were replaced by ${\mathcal{P}}$. In contrast, as currently stated, Theorem 3.6(ii) will only hold for ${\mathsf{P}}^{*}={\mathsf{Q}}$, which is the correct answer.
The situation improves in the presence of a reference measure $\mu$ with respect to which ${\mathsf{Q}}$ and all ${\mathsf{P}}\in{\mathcal{P}}$ are absolutely continuous. In this setting, if ${\mathsf{P}}_{n}$ and ${\mathsf{P}}$ are measures with densities $p_{n}=d{\mathsf{P}}_{n}/d\mu$ and $p=d{\mathsf{P}}/d\mu$ we say that ${\mathsf{P}}_{n}$ $\mu$-converges to ${\mathsf{P}}$ if $p_{n}$ converges to $p$ in probability under $\mu$. A set $C$ of absolutely continuous measures is called $\mu$-closed if it is closed with respect to $\mu$-convergence. It is called solid if it contains every (nonnegative) measure ${\mathsf{P}}^{\prime}$ that is dominated by some ${\mathsf{P}}\in C$ in the sense that ${\mathsf{P}}^{\prime}(A)\leq{\mathsf{P}}(A)$ for all $A$. The following result characterizes ${\mathcal{P}_{\textnormal{eff}}}$ in terms of these concepts.
@@
Suppose all ${\mathsf{P}}\in{\mathcal{P}}$ are absolutely continuous with respect to a probability measure $\mu$. Then all elements of ${\mathcal{P}_{\textnormal{eff}}}$ are also absolutely continuous with respect to $\mu$, and ${\mathcal{P}_{\textnormal{eff}}}$ is the smallest $\mu$-closed convex solid set that contains ${\mathcal{P}}$.
@@
We identify ${\mathcal{P}}$ with its set of densities $C=\{p\in L^{0}_{+}(\mu)\colon p=d{\mathsf{P}}/d\mu,\ {\mathsf{P}}\in{\mathcal{P}}\}$. The polar of $C$ in $L^{0}_{+}(\mu)$ is $C^{\circ}=\{f\in L^{0}_{+}(\mu)\colon{\mathbb{E}}_{\mu}[pf]\leq 1\text{ for all }p\in C\}$, which is the set of (equivalence classes of) random variables that are $\mu$-almost surely equal to a finite e-variable. We claim that all elements of ${\mathcal{P}_{\textnormal{eff}}}$ are absolutely continuous with respect to $\mu$, and that its set of densities is precisely the bipolar $C^{\circ\circ}=\{p\in L^{0}_{+}(\mu)\colon{\mathbb{E}}_{\mu}[pf]\leq 1\text{ for all }f\in C^{\circ}\}$. To check absolute continuity, let ${\mathsf{P}}\in{\mathcal{P}_{\textnormal{eff}}}$ and consider any event $A$ such that $\mu(A)=0$. Then $\infty\bm{1}_{A}$ is an e-variable, hence ${\mathbb{E}}_{\mathsf{P}}[\infty\bm{1}_{A}]\leq 1$, and thus ${\mathsf{P}}(A)=0$. This shows absolute continuity and lets us identify ${\mathcal{P}_{\textnormal{eff}}}$ with its set of densities, namely, the set of $p\in L^{0}_{+}(\mu)$ such that ${\mathbb{E}}_{\mu}[pX]\leq 1$ for all finite e-variables $X$. Here we may restrict to finite e-variables because ${\mathbb{E}}_{\mu}[pX]\leq 1$ if and only if ${\mathbb{E}}_{\mu}[p(X\wedge n)]\leq 1$ for all $n\in{\mathbb{N}}$. The finite e-variables are, up to $\mu$-nullsets, precisely the elements of $C^{\circ}$, so we deduce that ${\mathcal{P}_{\textnormal{eff}}}$ is indeed identified with $C^{\circ\circ}$. Finally, the bipolar theorem of Brannath and Schachermayer (1999, Theorem 1.3) states that $C^{\circ\circ}$ is the smallest $\mu$-closed convex solid set that contains $C$. This completes the proof. ∎
@@
If ${\mathcal{P}}$ is finite, then ${\mathcal{P}_{\textnormal{eff}}}$ consists of all sub-probabilities that are set-wise dominated by an element of $\operatorname{conv}({\mathcal{P}})$. In particular, ${\mathcal{P}_{\textnormal{eff}}}\cap M_{1}=\operatorname{conv}({\mathcal{P}})$.
@@
For finite ${\mathcal{P}}=\{{\mathsf{P}}_{1},\ldots,{\mathsf{P}}_{n}\}$ we may take $\mu=({\mathsf{P}}_{1}+\cdots+{\mathsf{P}}_{n})/n$ as reference measure. In this case $\operatorname{conv}({\mathcal{P}})$ is already $\mu$-closed and the form of ${\mathcal{P}_{\textnormal{eff}}}$ follows from Lemma 4.5. ∎
Lemma 4.5 is powerful because it characterizes all elements of ${\mathcal{P}_{\textnormal{eff}}}$. The following theorem makes use of this to simplify the conditions of Theorem 3.6. Lemma 4.5 also provides a way of checking that a particular measure belongs to ${\mathcal{P}_{\textnormal{eff}}}$. For example, it shows that limits in probability of (densities of) measures in the convex hull of ${\mathcal{P}}$ must belong to ${\mathcal{P}_{\textnormal{eff}}}$. The latter is however also easy to check directly from the definition of ${\mathcal{P}_{\textnormal{eff}}}$ and Fatou’s lemma, and does not require the bipolar theorem used in the proof of Lemma 4.5.
If a reference measure $\mu$ is given, we identify any ${\mathsf{P}}\in{\mathcal{P}}$ with its density $p=d{\mathsf{P}}/d\mu$. Thanks to Lemma 4.5, we may similarly identify ${\mathcal{P}_{\textnormal{eff}}}$ with its set of densities. For the sake of brevity we abuse notation and write, for example, $p\in{\mathcal{P}}$ even though, strictly speaking, ${\mathcal{P}}$ is a set of measures, not densities. Next, given ${\mathsf{Q}}$ with density $q$, the absolutely continuous part of any $p\in{\mathcal{P}_{\textnormal{eff}}}$ has density $p^{a}=p\bm{1}_{\{q>0\}}$, and this is ${\mathsf{Q}}$-almost surely equal to $p$ itself. In view of these observations, the conditions (ii) and (iii) of Theorem 3.6 characterizing the RIPr state, respectively, that ${\mathbb{E}}_{\mathsf{Q}}[p/p^{*}]\leq 1$ and ${\mathbb{E}}_{\mathsf{Q}}[\log(p/p^{*})]\leq 0$, for all $p\in{\mathcal{P}_{\textnormal{eff}}}$. Here $p^{*}=d{\mathsf{P}}^{*}/d\mu$. The following theorem simplifies these conditions by showing that, essentially, it is enough to check them for elements of ${\mathcal{P}}$ rather than ${\mathcal{P}_{\textnormal{eff}}}$.
@@
Assume that ${\mathsf{Q}}\ll{\mathcal{P}}$ and that ${\mathsf{Q}}$ and all elements of ${\mathcal{P}}$ are absolutely continuous with respect to a probability measure $\mu$. Let $p^{*}\in{\mathcal{P}_{\textnormal{eff}}}$ be equivalent to ${\mathsf{Q}}$. Then each of the conditions (i), (ii), and (iii) of Theorem 3.6 is equivalent to
@@
The reason for the asymmetry between Theorem 3.6(iii) and (4.4) is that the latter only involves $p$ from ${\mathcal{P}}$, not ${\mathcal{P}_{\textnormal{eff}}}$. Since ${\mathcal{P}}$ may not contain $p^{*}$, nor be convex, we need to explicitly include convex combinations with $p^{*}$ in the numerator on the left-hand side of (4.4). Let us also emphasize that $t=1$ is excluded in (4.4). This sometimes makes the condition easier to check, for instance in the proof of Corollary 4.10 below.
@@
We know that (i), (ii), and (iii) of Theorem 3.6 are equivalent for elements of ${\mathcal{P}_{\textnormal{eff}}}$ that are equivalent to ${\mathsf{Q}}$. To prove the theorem it is therefore enough to show that that Theorem 3.6(ii), (4.3), and (4.4) are equivalent.
We first show that Theorem 3.6(ii) is equivalent to (4.3). The former is clearly a stronger condition because it states that (4.3) holds with the larger set ${\mathcal{P}_{\textnormal{eff}}}$ in place of ${\mathcal{P}}$. For the converse, suppose (4.3) holds. Let ${\mathcal{P}}^{\prime}$ be the set of all sub-probability densities dominated by some element of the convex hull of ${\mathcal{P}}$, and let ${\mathcal{P}}^{\prime\prime}$ be the closure in $\mu$-probability of ${\mathcal{P}}^{\prime}$. It is clear that (4.3) holds for all $p\in{\mathcal{P}}^{\prime}$ and then, by Fatou’s lemma, also for all $p\in{\mathcal{P}}^{\prime\prime}$. We claim that ${\mathcal{P}}^{\prime\prime}$ is a closed convex solid set that contains ${\mathcal{P}}$; it must then also contain ${\mathcal{P}_{\textnormal{eff}}}$ since this is the smallest such set by Lemma 4.5. From this we conclude that Theorem 3.6(ii) holds. To see that ${\mathcal{P}}^{\prime\prime}$ is closed, convex, and solid, note that by definition, every $p^{\prime\prime}\in{\mathcal{P}}^{\prime\prime}$ is the limit in probability of some $p_{n}^{\prime}\in{\mathcal{P}}^{\prime}$, meaning that $p_{n}^{\prime}\leq p_{n}$ for some $p_{n}$ in the convex hull of ${\mathcal{P}}$. From this description convexity follows directly, closedness is obtained from a diagonal argument, and solidity follows by noting that any $\widetilde{p}\leq p^{\prime\prime}$ is the limit of $\widetilde{p}\wedge p_{n}^{\prime}$, which is still dominated by $p_{n}$ and thus belongs to ${\mathcal{P}}^{\prime}$.
@@
Even with a reference measure, it is possible that the relative entropy $H({\mathsf{Q}}\mid{\mathsf{P}})$ is infinite for all ${\mathsf{P}}\in{\mathcal{P}}$ but finite for the RIPr ${\mathsf{P}}^{*}$, similarly to what happens in Example 4.4. The following example illustrates this.
Example 4.9.
Let $\Omega={\mathbb{R}}$ with its Borel $\sigma$-algebra, let ${\mathcal{P}}$ be the set of all finite mixtures of Gaussians, and let ${\mathsf{Q}}$ be the standard Cauchy distribution. We may take $\mu={\mathsf{Q}}$ as the reference measure, for example. With this setup, $H({\mathsf{Q}}\mid{\mathsf{P}})=\infty$ for all ${\mathsf{P}}\in{\mathcal{P}}$. On the other hand, the Cauchy density can be obtained as the limit in $\mu$-probability (even in the uniform norm) of elements of ${\mathcal{P}}$, so it follows from Lemma 4.5 that ${\mathsf{Q}}\in{\mathcal{P}_{\textnormal{eff}}}$. Thus the RIPr is ${\mathsf{P}}^{*}={\mathsf{Q}}$ and $H({\mathsf{Q}}\mid{\mathsf{P}}^{*})=0$.
@@
if ${\mathsf{Q}}$ is absolutely continuous with respect to $p$, and +$\infty$+ ∞ otherwise. Note that this differs from the quantity $H({\mathsf{Q}}\mid{\mathsf{P}}\leadsto{\mathcal{P}_{\textnormal{eff}}})$ in Remark 3.9, which involves ${\mathcal{P}_{\textnormal{eff}}}$ rather than ${\mathcal{P}}$.
@@
Assume that ${\mathsf{Q}}\ll{\mathcal{P}}$ and that all elements of ${\mathcal{P}}$ are absolutely continuous with respect to a probability measure $\mu$. If ${\mathcal{P}}$ is convex and $(p_{n})_{n\in{\mathbb{N}}}$ is a sequence in ${\mathcal{P}}$ with $H({\mathsf{Q}}\mid p_{n}\leadsto{\mathcal{P}})\to 0$, then $\log p_{n}$ converges to $\log p^{*}$ in $L^{1}({\mathsf{Q}})$, where $p^{*}$ is the RIPr.
@@
Theorem 4.7 now shows that $p^{*}$ is the RIPr, provided we can argue that $p^{*}$ belongs to ${\mathcal{P}_{\textnormal{eff}}}$. This will use that $p^{*}=0$ on $\{q=0\}$. Specifically, note that ${\mathcal{P}_{\textnormal{eff}}}$ contains $p_{n}$ and hence, being solid, also the smaller random variable $p_{n}\bm{1}_{\{q>0\}}$. Convergence in ${\mathsf{Q}}$-probability of $p_{n}$ to $p^{*}$ implies convergence in $\mu$-probability of $p_{n}\bm{1}_{\{q>0\}}$ to $p^{*}\bm{1}_{\{q>0\}}=p^{*}$. Since ${\mathcal{P}_{\textnormal{eff}}}$ is closed in $\mu$-probability, it must contain $p^{*}$. This completes the proof. ∎